OpenStudy (anonymous):

Stumped on a question.. (-1/x )+( 2/x^2+1 )+ ( 1/x^3+x ) Thank you! :)

OpenStudy (anonymous):

least common multiple of the denominators is \[x^3+x\] so multiply first one top and bottom by \[x^2+1\] and the second one top and bottom by x and third one leave alone. then add in the numerator

OpenStudy (anonymous):

\[\frac{-1}{x}\frac{(x^2+1)}{(x^2+1)}+\frac{2}{(x^2+1)}\frac{x}{x}+\frac{1}{x^3+x}\]

OpenStudy (anonymous):

get \[\frac{-x^2-1+2x+1}{x^3+x}\] \[=\frac{-x^2+2x}{x^3+x}\] \[=\frac{x(2-x)}{x(x^2+1)}\] \[=\frac{2-x}{x^2+1}\]

OpenStudy (anonymous):

Oh woah, the way you explained the question made it so simple. Thanks a lot!